# MDNS: Masked Diffusion Neural Sampler via Stochastic Optimal Control

### v1

@@ Abstract
We study the problem of learning a neural sampler to generate samples from discrete state spaces where the target probability mass function $\pi\propto\mathrm{e}^{-U}$ is known up to a normalizing constant, which is an important task in fields such as statistical physics, machine learning, combinatorial optimization, etc. To better address this challenging task when the state space has a large cardinality and the distribution is multi-modal, we propose Masked Diffusion Neural Sampler (MDNS), a novel framework for training discrete neural samplers by aligning two path measures through a family of learning objectives, theoretically grounded in the stochastic optimal control of the continuous-time Markov chains. We validate the efficiency and scalability of MDNS through extensive experiments on various distributions with distinct statistical properties, where MDNS learns to accurately sample from the target distributions despite the extremely high problem dimensions and outperforms other learning-based baselines by a large margin. A comprehensive study of ablations and extensions is also provided to demonstrate the efficacy and potential of the proposed framework.
@@ 1 Introduction
Drawing samples from an unnormalized target distribution $\pi\propto\mathrm{e}^{-U}$ on some state space $\mathcal{X}_{0}$ is a fundamental problem with wide-ranging applications across fields such as statistical physics [LB14], Bayesian inference [Gel+13], machine learning [And+03, WJ08], etc. For decades, Markov chain Monte Carlo (MCMC) methods, such as the Langevin Monte Carlo, the Metropolis-Hastings algorithm, and the Glauber dynamics, have been a cornerstone. These methods simulate a Markov chain whose stationary distribution is the target distribution, and converge provably fast under certain circumstances [Che22]. Despite their widespread adoption, MCMC algorithms face significant challenges, especially when the state space has high dimensions and the target distribution is multimodal, which hinders efficient exploration and convergence (see, e.g., the lower bound results in [Ran06, GT06, GLR18, HZ25]).
@@
Recently, inspired by the success of diffusion models in generative modeling for continuous data [HJA20, SME21, Son+21], significant progress has been made in leveraging neural networks to learn a stochastic differential equation (SDE) to generate samples from $\pi$, which we collectively refer to as neural samplers (e.g., [ZC22, VGD23, AVE25, Hav+25]). Concurrently, the diffusion paradigm has been effectively extended to discrete state spaces [Aus+21, Cam+22, Sun+23a, LME24], finding broad applications in domains involving sequences of categorical data. However, while neural samplers have been extensively developed for distributions on $\mathbb{R}^{d}$ and discrete diffusion models have shown promise in modeling discrete data, the study of leveraging discrete diffusion models for sampling remains relatively unexplored. Addressing this gap is crucial as numerous real-world sampling problems in areas like statistical physics [FL24] and combinatorial optimization [Sun+23] necessitate specialized methods to navigate the unique challenges they present, and this work aims to develop such a specialized approach.
Contributions. Our contributions can be summarized in the following points. (1) We introduce Masked Diffusion Neural Sampler (MDNS), a novel framework for training discrete neural samplers leveraging optimal control of CTMC and masked discrete diffusion models. (2) To cope with the optimization challenges caused by the discontinuity of CTMC trajectories, we introduce several learning objectives that are free of differentiability requirements. (3) To achieve scalable training in high dimensions, we propose a novel training loss named Weighted Denoising Cross-entropy ˜16 that directly implements score learning on the target distribution using importance sampling. (4) We conduct comprehensive experiments and ablation studies on Ising and Potts models to validate the efficacy of our method. MDNS can accurately sample from target distributions even when the state space has a cardinality of $10^{122}$, which substantially surpasses other learning-based methods.
@@ 2.2 Discrete Diffusion Models
In discrete diffusion models, one learns the generator of a CTMC that starts from a simple distribution $p_{\mathrm{init}}$ and reaches the target distribution $p_{\mathrm{data}}$ at the final time. This CTMC is typically chosen as the time-reversal of a noising process that converts any distribution to $p_{\mathrm{init}}$. An especially effective subclass of discrete diffusion is the masked discrete diffusion models [LME24, Ou+25, Sah+24, Shi+24, Zhe+25], which corresponds to choosing $p_{\mathrm{init}}$ to be $p_{\mathrm{mask}}$, the Dirac distribution on the fully masked sequence. Adding a mask token $\mathbf{M}$ into the original state space $\mathcal{X}_{0}:=\{1,2,...,N\}^{D}$ containing length-$N$ sequences with vocabulary size $D$, denote the mask-augmented state space by $\mathcal{X}:=\{1,2,...,N,\mathbf{M}\}^{D}$. Throughout the paper, $x=(x^{1},...,x^{D})$ denotes a sequence of tokens, $x^{d\leftarrow n}$ represents the sequence constructed by replacing the $d$-th position of $x$ by $n$, and $x^{\mathrm{UM}}=(x^{d}:x^{d}\neq\mathbf{M})$ represents the unmasked part of $x\in\mathcal{X}$.
@@
and 0 if otherwise, for some noise schedule $\gamma:[0,T]\to[0,\infty)$. In practice, practitioners typically leverage a neural network $s_{\theta}$ that takes a partially masked sequence $x\in\mathcal{X}$ as input and outputs $s_{\theta}(x)\in\mathbb{R}^{D\times N}$ whose $(d,n)$ entry approximates $\Pr_{X\sim p_{\mathrm{data}}}(X^{d}=n|X^{\mathrm{UM}}=x^{\mathrm{UM}})$ if $x^{d}=\mathbf{M}$.
@@ 3.2 Optimal Control Formulation of Discrete Neural Sampler Training
It is straightforward to see that the problem ˜6 is a special case of the proposed framework upon choosing $\mathcal{F}(\mathbb{P}^{u},\mathbb{P}^{*})=\operatorname{KL}(\mathbb{P}^{u}\|\mathbb{P}^{*})$. However, while being directly relevant to many theoretical results, naively optimizing the discretized, estimated objective in ˜6 is not an appropriate approach to learn the desired controlled generator $Q^{u}$. The estimation of ˜6 requires the simulation of CTMC using a neural network parameterized generator with parameter $\theta$, yet the objective itself is inherently non-differentiable with respect to $\theta$ since (1) the trajectories of CTMCs are pure jump processes, and (2) the reward function $r$ is non-differentiable, making it difficult to effectively train the generator.
A workaround is proposed in [Wan+25], where the authors retain the differentiability of the objectives through relaxing the CTMC trajectories from staying on $\mathcal{X}_{0}$ to sequences of probability vectors on $\mathbb{R}^{D\times N}$, using the Gumbel softmax trick. This partly solves the problem at the cost of introducing a high approximation error into the learning process, since the $\theta$ gradients coming from the Gumbel softmax trick are known to be biased and easily cause numerical instability [JGP17, Liu+23]. This causes a failure to converge to the correct target distribution even in low dimensions, as is validated empirically in our experiments (see Sec.˜D.5). Moreover, the backpropagation through the whole trajectory is a memory-intensive operation, and we would like to avoid it as much as possible. In the following, we propose several alternative learning objectives that operate without these disadvantages.
Relative-entropy with REINFORCE (RERF). One of the key reasons that $\operatorname{KL}(\mathbb{P}^{u}\|\mathbb{P}^{*})$ is intractable for direct optimization originates from the fact that the expectation is with respect to $\mathbb{P}^{u}$, which tracks the $\theta$ gradient. Alternatively, we can estimate $\nabla_{\theta}\operatorname{KL}(\mathbb{P}^{u}\|\mathbb{P}^{*})$ by introducing a gradient-nontracking trajectory simulated using ${\bar{u}}=\texttt{stopgrad}\left(u\right)$, also known as the REINFORCE trick [Wil92, RGB14, MG14]. We denote the path measure produced with $\bar{u}$} to be $\mathbb{P}^{{\bar{u}}}$, [^2] This means we sample trajectory $X\sim\mathbb{P}^{u}$ and detach it from the computational graph. After that, use the values of the trajectory to compute the objective and do backpropagationf. and express the gradient of the relative-entropy by the following identity: $\nabla_{\theta}\operatorname{KL}(\mathbb{P}^{u}\|\mathbb{P}^{*})=\nabla_{\theta}\mathcal{F}_{\mathrm{RERF}}(\mathbb{P}^{u},\mathbb{P}^{*})$, where
@@
See Sec.˜C.3 for a detailed derivation. Thus, $\mathcal{F}_{\mathrm{RERF}}$ can be introduced for the purpose of gradient estimation of relative-entropy. We remark that in contrast to the approach proposed in [Wan+25], ˜9 gives an unbiased estimator of the relative-entropy gradient, which is crucial for the accurate learning of the target distribution. However, $\mathcal{F}_{\mathrm{RERF}}(\mathbb{P}^{u},\mathbb{P}^{*})$ is not a valid “loss function” but only a computational surrogate, in the sense that a reduction in the objective value does not guarantee an improvement in learning performance.
@@ 3.3 Masked Diffusion Neural Sampler
Besides the objective $\mathcal{F}(\mathbb{P}^{u},\mathbb{P}^{*})$, another major component in the design space of the proposed training framework is the choice of reference path measure $\mathbb{P}^{0}$ and the corresponding generator $Q^{0}$. Recall from Sec.˜3.1 that we need the reference path measure to be memoryless to guarantee the existence and uniqueness of the SOC problem ˜6. We now introduce a method for choosing such memoryless reference path measure based on a masked discrete diffusion model. This approach, which we term the Masked Diffusion Neural Sampler (MDNS), forms the core of our framework. The corresponding learning algorithms are subsequently developed based on the objectives proposed in Sec.˜3.2. Furthermore, we demonstrate that this framework can be extended to incorporate uniform discrete diffusion model, an adaptation referred to as the Uniform Diffusion Neural Sampler (UDNS), with a detailed discussion deferred to App.˜F.
@@
To guarantee the reference path measure $\mathbb{P}^{0}$ with generator $Q^{0}$ defined in ˜12 and starting from $p_{\mathrm{mask}}$ satisfies $\mathbb{P}^{0}_{T}=p_{\mathrm{unif}}$ and is memoryless, it suffices to ensure $\int_{0}^{T}\gamma(t)\mathrm{d}t=\infty$.
@@ 4 Experiments
In this section, we experimentally validate our proposed frameworks by learning to sample from the Ising model and Potts model on square lattices. At different temperatures, these models exhibit distinct behaviors, which provide a rich ground to demonstrate the effectiveness of our algorithms. We use probabilistic metrics (KL divergence, TV distance, $\chi^{2}$ divergence, etc.) and observables (magnetization, 2-point correlation) originating from statistical physics to benchmark our learning results. We include both learning-based approaches such as LEAPS [HAJ25], and learning-free approaches such as the Metropolis-Hastings (MH) algorithm as benchmarks. For fair comparisons, we run all baselines with a long enough but comparable compute time. To accurately estimate ground truth values for all the observables when an exact computation is intractable, we draw samples by running the Swendsen-Wang (SW) algorithm [SW86, SW87] for a sufficiently long time. We include experiment details such as metric definitions and training hyperparameters in Apps.˜D and E.
@@ 4.1 Ising model on Square Lattice
A case study on the choice of $\mathcal{F}$. We first compare the effects of different learning objectives $\mathcal{F}$ by learning to sample from a $4\times 4$ high temperature Ising model with $J=1$, $h=0.1$, and $\beta_{\mathrm{high}}=0.28$. The system size is picked so that the sampling problem is not trivial, but an exact computation of the partition function remains computationally tractable. The results are reported in Tab.˜2, and results at other temperatures can be found at App.˜D. For a fair comparison, we train for $1000$ steps with batch size $256$ for each loss and pick the resampling replicates $R=16$ for $\mathcal{F}_{\mathrm{WDCE}}$ so that the effective batch size is the same for all objectives. We draw a sufficiently large number ($2^{20}$) of samples from the learned model to estimate all the reported metrics. We also report the average effective sample size (ESS, defined in ˜24) of the last $100$ steps to showcase the learning speed for each objective. We include a baseline produced by running the MH algorithm for sufficiently long as a reference. We can see that all four learning objectives are able to learn a good model that generates ground truth-like samples. In all the evaluation metrics except the absolute error of $\log\widehat{Z}$, the effectiveness ranking is uniformly $\mathcal{F}_{\mathrm{LV}}>\mathcal{F}_{\mathrm{WDCE}}>\mathcal{F}_{\mathrm{RERF}}>\mathcal{F}_{\mathrm{CE}}$.
@@
Table 1: Results for learning $16\times 16$ Ising model at different temperatures, best in bold. Mag. and Corr. represent absolute error of the magnetization and 2-point correlation to ground truth values.
@@
Scaling to higher dimensions and lower temperatures. We also run on significantly harder settings by increasing the system size to $16\times 16$ (so the cardinality of the state space is $2^{256}$). We keep $J=1,h=0$ and vary the inverse temperature, ranging from $\beta_{\mathrm{high}}=0.28$, $\beta_{\mathrm{critical}}=\log(1+\sqrt{2})/2=0.4407$, to $\beta_{\mathrm{low}}=0.6$. The behaviors of the Ising model are known to be distinct at these temperatures by the phase transition theory [KW41, Ons44]. For the purpose of efficient training, we use $\mathcal{F}_{\mathrm{WDCE}}$ as learning objectives and train on each problem for $50\mathrm{k}$ steps. In Tab.˜1, we report the ESS and the absolute error of magnetization and 2-point correlation (respectively defined in ˜26 and 28) to the estimated ground truth. We also plot the estimated average 2-point correlation (defined in (27)) between a varying distance $r$ in Fig.˜1, and defer further results to Sec.˜D.3. Our learned distribution accurately matches the theoretically predicted trend of correlation functions, which exponentially decays to the non-zero spontaneous magnetization at $\beta_{\mathrm{low}}$, polynomial decays at $\beta_{\mathrm{critical}}$ and exponential decays to 0 at $\beta_{\mathrm{high}}$. As is evident from the table and the figure, MDNS outperforms LEAPS by a large margin in almost all the metrics. MDNS accurately learns the high-dimensional distribution across all temperatures suggested by the low error in observables and high ESS values, while LEAPS fails at $\beta_{\mathrm{critical}}$ and $\beta_{\mathrm{low}}$ despite having been trained for a long time.
Warm-up for lower temperatures. Instead of only training from scratch, our training includes a warm-up phase. When targeting hard distributions such as $\beta_{\mathrm{critical}}$ and $\beta_{\mathrm{low}}$, to help the model better locate the modes of the target distribution, we start the training by fitting easier ones such as $\beta_{\mathrm{high}}$ for a short ($20\mathrm{k}$ steps in this example) period of time. An ablation study to demonstrate the importance of this proposed technique is presented in Sec.˜D.2.

### v2

@@ Abstract
We study the problem of learning a neural sampler to generate samples from discrete state spaces where the target probability mass function $\pi\propto\mathrm{e}^{-U}$ is known up to a normalizing constant, which is an important task in fields such as statistical physics, machine learning, combinatorial optimization, etc. To better address this challenging task when the state space has a large cardinality and the distribution is multi-modal, we propose Masked Diffusion Neural Sampler (MDNS), a novel framework for training discrete neural samplers by aligning two path measures through a family of learning objectives, theoretically grounded in the stochastic optimal control of the continuous-time Markov chains. We validate the efficiency and scalability of MDNS through extensive experiments on various distributions with distinct statistical properties, where MDNS learns to accurately sample from the target distributions despite the extremely high problem dimensions and outperforms other learning-based baselines by a large margin. A comprehensive study of ablations and extensions is also provided to demonstrate the efficacy and potential of the proposed framework. Our code is available at https://github.com/yuchen-zhu-zyc/MDNS.
@@ 1 Introduction
Drawing samples from an unnormalized target distribution $\pi\propto\mathrm{e}^{-U}$ on some state space $\mathcal{X}_{0}$ is a fundamental problem with wide-ranging applications across fields such as statistical physics [LB14], Bayesian inference [Gel+13], machine learning [And+03, WJ08], etc. For decades, Markov chain Monte Carlo (MCMC) methods, such as the Langevin Monte Carlo, the Metropolis-Hastings algorithm, and the Glauber dynamics, have been a cornerstone. These methods simulate a Markov chain whose stationary distribution is the target distribution, and converge provably fast under certain circumstances [Che22]. Despite their widespread adoption, MCMC algorithms face significant challenges, especially when the state space is high-dimensional and the target distribution is multimodal, which hinders efficient exploration and convergence (see, e.g., the lower-bound results in [Ran06, GT06, GLR18, HZ25]).
@@
Recently, inspired by the success of diffusion models in generative modeling for continuous data [HJA20, SME21, Son+21], significant progress has been made in leveraging neural networks to learn a stochastic differential equation (SDE) to generate samples from $\pi$, which we collectively refer to as neural samplers (e.g., [ZC22, VGD23, AVE25, Hav+25]). Concurrently, the diffusion paradigm has been effectively extended to discrete state spaces [Aus+21, Cam+22, Sun+23a, LME24], finding broad applications in domains involving sequences of categorical data. However, while neural samplers have been extensively developed for distributions on $\mathbb{R}^{d}$ and discrete diffusion models have shown promise in modeling discrete data, the study of leveraging discrete diffusion models for sampling remains relatively unexplored. Addressing this gap is crucial, as numerous real-world sampling problems in areas such as statistical physics [FL24] and combinatorial optimization [Sun+23, SHL24, San+25] necessitate specialized methods to navigate the unique challenges they pose. This work aims to develop such a specialized approach.
Contributions. Our contributions can be summarized in the following points.
  1. 1.
    We introduce Masked Diffusion Neural Sampler (MDNS), a novel framework for training discrete neural samplers that leverages optimal control of CTMCs and masked discrete diffusion models.
  2. 2.
    To cope with the optimization challenges caused by the discontinuity of CTMC trajectories, we introduce several learning objectives that are free of differentiability requirements.
  3. 3.
    To achieve scalable training in high dimensions, we propose a novel training loss named Weighted Denoising Cross-entropy ˜16 that directly implements score learning on the target distribution using importance sampling.
  4. 4.
    We conduct comprehensive experiments and ablation studies on Ising and Potts models to validate the efficacy of our method. MDNS can accurately sample from target distributions even when the state space has cardinality $10^{122}$, substantially surpassing other learning-based methods.
@@ 2.2 Discrete Diffusion Models
In discrete diffusion models, one learns the generator of a CTMC that starts from a simple distribution $p_{\mathrm{init}}$ and reaches the target distribution $p_{\mathrm{data}}$ at the final time. This CTMC is typically chosen as the time-reversal of a noising process that converts any distribution to $p_{\mathrm{init}}$. An especially effective subclass of discrete diffusion is the masked discrete diffusion models [LME24, Ou+25, Sah+24, Shi+24, Zhe+25], which corresponds to choosing $p_{\mathrm{init}}$ to be $p_{\mathrm{mask}}$, the Dirac distribution on the fully masked sequence. Adding a mask token $\mathbf{M}$ into the original state space $\mathcal{X}_{0}:=\{1,2,...,N\}^{D}$ containing length-$D$ sequences with vocabulary size $N$, denote the mask-augmented state space by $\mathcal{X}:=\{1,2,...,N,\mathbf{M}\}^{D}$. Throughout the paper, $x=(x^{1},...,x^{D})$ denotes a sequence of tokens, $x^{d\leftarrow n}$ represents the sequence constructed by replacing the $d$-th position of $x$ by $n$, and $x^{\mathrm{UM}}=(x^{d}:x^{d}\neq\mathbf{M})$ represents the unmasked part of $x\in\mathcal{X}$.
@@
and 0 if otherwise, for some noise schedule $\gamma:[0,T]\to[0,\infty)$ such that $\int_{0}^{T}\gamma(t)\mathrm{d}t=\infty$. In practice, practitioners typically leverage a neural network $s_{\theta}$ that takes a partially masked sequence $x\in\mathcal{X}$ as input and outputs $s_{\theta}(x)\in\mathbb{R}^{D\times N}$ whose $(d,n)$ entry approximates $\Pr_{X\sim p_{\mathrm{data}}}(X^{d}=n|X^{\mathrm{UM}}=x^{\mathrm{UM}})$ if $x^{d}=\mathbf{M}$.
@@ 3.2 Optimal Control Formulation of Discrete Neural Sampler Training
It is straightforward to see that the problem ˜6 is a special case of the proposed framework upon choosing $\mathcal{F}(\mathbb{P}^{u},\mathbb{P}^{*})=\operatorname{KL}(\mathbb{P}^{u}\|\mathbb{P}^{*})$. However, although directly relevant to many theoretical results, naively optimizing the discretized, estimated objective in ˜6 is not an appropriate approach for learning the desired controlled generator $Q^{u}$. The estimation of ˜6 requires the simulation of CTMC using a neural network parameterized generator with parameter $\theta$, yet the objective itself is inherently non-differentiable with respect to $\theta$ since (1) the trajectories of CTMCs are pure jump processes, and (2) the reward function $r$ is non-differentiable, making it difficult to effectively train the generator.
A workaround is proposed in [Wan+25], where the authors retain the differentiability of the objectives through relaxing the CTMC trajectories from staying on $\mathcal{X}_{0}$ to sequences of probability vectors on $\mathbb{R}^{D\times N}$, using the Gumbel softmax trick. This partly solves the problem at the cost of introducing a high approximation error into the learning process, since the $\theta$ gradients coming from the Gumbel softmax trick are known to be biased and easily cause numerical instability [JGP17, Liu+23]. This leads to failure to converge to the correct target distribution, even in low dimensions, as empirically validated in our experiments (see Sec.˜D.5). Moreover, backpropagation over the entire trajectory is a memory-intensive operation, and we would like to avoid it as much as possible. In the following, we propose several alternative learning objectives that operate without these disadvantages.
Relative-entropy with REINFORCE (RERF). One of the key reasons that $\operatorname{KL}(\mathbb{P}^{u}\|\mathbb{P}^{*})$ is intractable for direct optimization originates from the fact that the expectation is with respect to $\mathbb{P}^{u}$, which tracks the $\theta$ gradient. Alternatively, we can estimate $\nabla_{\theta}\operatorname{KL}(\mathbb{P}^{u}\|\mathbb{P}^{*})$ by introducing a gradient-nontracking trajectory simulated using ${\bar{u}}=\texttt{stopgrad}\left(u\right)$, also known as the REINFORCE trick [Wil92, RGB14, MG14]. We denote the path measure produced with $\bar{u}$} to be $\mathbb{P}^{{\bar{u}}}$ [^2] This means we sample a trajectory $X\sim\mathbb{P}^{u}$ and detach it from the computational graph. After that, use the values of the trajectory to compute the objective and do backpropagation., and express the gradient of the relative-entropy by the following identity: $\nabla_{\theta}\operatorname{KL}(\mathbb{P}^{u}\|\mathbb{P}^{*})=\nabla_{\theta}\mathcal{F}_{\mathrm{RERF}}(\mathbb{P}^{u},\mathbb{P}^{*})$, where
@@
See Sec.˜C.3 for a detailed derivation. Thus, $\mathcal{F}_{\mathrm{RERF}}$ can be introduced for the purpose of gradient estimation of relative entropy. We remark that, in contrast to the approach proposed in [Wan+25], ˜9 yields an unbiased estimator of the relative-entropy gradient, which is crucial for accurate learning of the target distribution. However, $\mathcal{F}_{\mathrm{RERF}}(\mathbb{P}^{u},\mathbb{P}^{*})$ is not a valid “loss function” but only a computational surrogate, in the sense that a reduction in the objective value does not guarantee an improvement in learning performance.
@@ 3.3 Masked Diffusion Neural Sampler
Besides the objective $\mathcal{F}(\mathbb{P}^{u},\mathbb{P}^{*})$, another major component in the design space of the proposed training framework is the choice of reference path measure $\mathbb{P}^{0}$ and the corresponding generator $Q^{0}$. Recall from Sec.˜3.1 that we need the reference path measure to be memoryless to guarantee the existence and uniqueness of the SOC problem ˜6. We now introduce a method for choosing such a memoryless reference path measure based on a masked discrete diffusion model. This approach, which we term the Masked Diffusion Neural Sampler (MDNS), forms the core of our framework. The corresponding learning algorithms are subsequently developed based on the objectives proposed in Sec.˜3.2. Furthermore, we demonstrate that this framework can be extended to incorporate uniform discrete diffusion models, an adaptation referred to as the Uniform Diffusion Neural Sampler (UDNS), with a detailed discussion deferred to App.˜F.
@@
Under the assumption that $\int_{0}^{T}\gamma(t)\mathrm{d}t=\infty$, the reference path measure $\mathbb{P}^{0}$ with generator $Q^{0}$ defined in ˜12 and starting from $p_{\mathrm{mask}}$ is memoryless and satisfies $\mathbb{P}^{0}_{T}=p_{\mathrm{unif}}$.
@@ 4 Experiments
In this section, we experimentally validate our proposed frameworks by learning to sample from the Ising model and Potts model on square lattices. At different temperatures, these models exhibit distinct behaviors, providing a rich ground for demonstrating the effectiveness of our algorithms. We use probabilistic metrics (KL divergence, TV distance, $\chi^{2}$ divergence, etc.) and observables (magnetization, 2-point correlation) originating from statistical physics to benchmark our learning results. We include both learning-based approaches such as LEAPS [HAJ25], and learning-free approaches such as the Metropolis-Hastings (MH) algorithm as benchmarks. For fair comparisons, we run all baselines with a long enough but comparable compute time. To accurately estimate ground truth values for all the observables when an exact computation is intractable, we draw samples by running the Swendsen-Wang (SW) algorithm [SW86, SW87] for a sufficiently long time. We include experiment details such as metric definitions and training hyperparameters in Apps.˜D and E.
@@ 4.1 Ising model on Square Lattice
A case study on the choice of $\mathcal{F}$. We first compare the effects of different learning objectives $\mathcal{F}$ by learning to sample from a $4\times 4$ high temperature Ising model with $J=1$, $h=0.1$, and $\beta_{\mathrm{high}}=0.28$. The system size is picked so that the sampling problem is not trivial, but an exact computation of the partition function remains computationally tractable. The results are reported in Tab.˜2, and results at other temperatures are available in App.˜D. For a fair comparison, we train for $1000$ steps with a batch size of $256$ for each loss, and use $R=16$ resampling replicates for $\mathcal{F}_{\mathrm{WDCE}}$ so that the effective batch size is the same across all objectives. We draw a sufficiently large number ($2^{20}$) of samples from the learned model to estimate all the reported metrics. We also report the average effective sample size (ESS, defined in ˜24) of the last $100$ steps to showcase the learning speed for each objective. We include a baseline produced by running the MH algorithm for sufficiently long as a reference. We can see that all four learning objectives can learn a good model that generates ground-truth-like samples. In all the evaluation metrics except the absolute error of $\log\widehat{Z}$, the effectiveness ranking is uniformly $\mathcal{F}_{\mathrm{LV}}>\mathcal{F}_{\mathrm{WDCE}}>\mathcal{F}_{\mathrm{RERF}}>\mathcal{F}_{\mathrm{CE}}$.
@@
Table 1: Results for learning $16\times 16$ Ising model at different temperatures, best in bold. Mag. and Corr. represent the absolute error of the magnetization and 2-point correlation to ground truth values.
@@
Scaling to higher dimensions and lower temperatures. We also run on significantly harder settings by increasing the system size to $16\times 16$ (so the cardinality of the state space is $2^{256}$). We keep $J=1,h=0$ and vary the inverse temperature, ranging from $\beta_{\mathrm{high}}=0.28$, $\beta_{\mathrm{critical}}=\log(1+\sqrt{2})/2=0.4407$, to $\beta_{\mathrm{low}}=0.6$. The behaviors of the Ising model are known to be distinct at these temperatures by the phase transition theory [KW41, Ons44]. For the purpose of efficient training, we use $\mathcal{F}_{\mathrm{WDCE}}$ as learning objectives and train on each problem for $50\mathrm{k}$ steps. In Tab.˜1, we report the ESS and the absolute error of magnetization and 2-point correlation (respectively defined in ˜26 and 28) to the estimated ground truth. We also plot the estimated average 2-point correlation (defined in (27)) between a varying distance $r$ in Fig.˜1, and defer further results to Sec.˜D.3. Our learned distribution accurately matches the theoretically predicted trend of correlation functions, which exponentially decays to the non-zero spontaneous magnetization at $\beta_{\mathrm{low}}$, polynomial decays at $\beta_{\mathrm{critical}}$ and exponential decays to 0 at $\beta_{\mathrm{high}}$. As shown in the table and figure, MDNS outperforms LEAPS by a large margin across almost all metrics. MDNS accurately learns the high-dimensional distribution across all temperatures, as evidenced by low observables errors and high ESS values, whereas LEAPS fails at $\beta_{\mathrm{critical}}$ and $\beta_{\mathrm{low}}$ despite long training.
Warm-up for lower temperatures. Instead of only training from scratch, our training includes a warm-up phase. When targeting hard distributions such as $\beta_{\mathrm{critical}}$ and $\beta_{\mathrm{low}}$, to help the model better locate the modes of the target distribution, we start the training by fitting easier ones such as $\beta_{\mathrm{high}}$ for a short ($20\mathrm{k}$ steps in this example) period of time. An ablation study to demonstrate the importance of this proposed technique is presented in Sec.˜D.2. A more systematic study on the warm-up is deferred to future work [Guo+25].